I need help on the problem: 5(-1)-2(2(-1)-8)
Please and Thank You!!!

Answer by Maricel Balos(8)

5(-1)-2[2(-1)-8]
=-5-2[-2-8]
=-5-2(-10)
=-5-(-20)
=-5+20
=15

Therefore the ans. is 15